OpenStudy (anonymous):

im having trouble with Vector Valued Functions. How do I find the domain of a fn? ex: r(t)=5ti-4tj-(1/t)k

OpenStudy (anonymous):

take each component by itself and find the domain of each

OpenStudy (amistre64):

t cant be 0

OpenStudy (anonymous):

5t can be any real number

OpenStudy (anonymous):

-4t can be any real number however the last 1/t will become undefined at t=0

OpenStudy (anonymous):

so the domain has to fit all the components so for what values can all of these components get a value

OpenStudy (anonymous):

so its all real numbers not including 0?

OpenStudy (anonymous):

the answer would be all real numbers EXCEPT for 0 or \[(-\infty,0)\cup(0,\infty)\]

OpenStudy (anonymous):

awesome! that's what my book says and now i understand why. thank you!